OpenStudy (anonymous):

change -54 degrees to radian measure in terms of pi

OpenStudy (johnweldon1993):

To convert from degrees to radians you multiply by pi/180 \[\large -54 \times \frac{\pi}{180}\] and then simplify

OpenStudy (anonymous):

i tried -54(pi/180 and got 1?

OpenStudy (anonymous):

-0.942477796076938?

OpenStudy (shamil98):

in terms of pi, means you dont convert it to 3.14...

OpenStudy (johnweldon1993):

So like this... \[\large -54 \times \frac{\pi}{180} = \frac{-54\pi}{180}\] You would just simplify...not evaluate

OpenStudy (shamil98):

you reduce what you can in the numerator and denominator (the constants of (-54pi/180)

OpenStudy (anonymous):

oh ok so just reduce 54 and 180?

OpenStudy (sweetburger):

indeed

OpenStudy (johnweldon1993):

Right...so what would it reduce to?

OpenStudy (jdoe0001):

\(\bf radians = \cfrac{degrees\cdot \pi}{180}\)

OpenStudy (anonymous):

-5pi/12?

OpenStudy (jdoe0001):

\(\bf \cfrac{-54}{180}\ne \cfrac{5}{12}\)

OpenStudy (anonymous):

-3pi/10?

OpenStudy (johnweldon1993):

There we go!

OpenStudy (jdoe0001):

\(\bf \cfrac{-54}{180}\ne \cfrac{-5}{12}\) rather

OpenStudy (anonymous):

thank you not sure who to give the medal you all helped.